# Multi-Agent Collaborative Framework for Intelligent IT Operations: An AOI System with Context-Aware Compression and Dynamic Task Scheduling

### v1

@@ Abstract
The proliferation of cloud-native architectures, characterized by microservices and dynamic orchestration, has rendered modern IT infrastructures exceedingly complex and volatile. This complexity generates overwhelming volumes of operational data, leading to critical bottlenecks in conventional systems: inefficient information processing, poor task coordination, and loss of contextual continuity during fault diagnosis and remediation. To address these challenges, we propose AOI (AI-Oriented Operations), a novel multi-agent collaborative framework that integrates three specialized agents with an LLM-based Context Compressor. Its core innovations include: (1) a dynamic task scheduling strategy that adaptively prioritizes operations based on real-time system states, and (2) a three-layer memory architecture comprising Working, Episodic, and Semantic layers that optimizes context retention and retrieval. Extensive experiments on both synthetic and real-world benchmarks demonstrate that AOI effectively mitigates information overload—achieving a 72.4% context compression ratio while preserving 92.8% of critical information—and significantly enhances operational efficiency, attaining a 94.2% task success rate and reducing the Mean Time to Repair (MTTR) by 34.4% compared to the best baseline. This work presents a paradigm shift towards scalable, adaptive, and context-aware autonomous operations, enabling robust management of next-generation IT infrastructures with minimal human intervention.
I Introduction
@@
Modern IT infrastructures have evolved into highly distributed, cloud-native systems composed of microservices, containerized workloads, and dynamically orchestrated computing resources [dragoni2017microservices]. While this architectural paradigm significantly enhances scalability and deployment agility, it also introduces unprecedented complexity and volatility into system behaviors [gannon2017cloud]. This shift not only surpasses the cognitive limits of human operators but also fundamentally challenges the design principles of automated operation systems, which must now reason about highly dynamic, interdependent, and large-scale system states [joy2024aiops]. As organizations scale their digital services, the volume of logs, monitoring metrics, and alerts grows exponentially, quickly surpassing the cognitive and operational limits of human operators [brewer2016sre]. This challenge is well illustrated by large-scale production environments, where even a single misconfiguration or latent performance anomaly can propagate unpredictably across microservices, complicating diagnosis and dramatically increasing mean time to recovery (MTTR) [taibi2019microservices]. Empirical studies show that log data in modern distributed systems can reach terabyte-level daily volumes, making timely incident detection and remediation increasingly infeasible without automation [xu2009detecting]. Consequently, the demand for intelligent, autonomous IT operations systems has become critical for sustaining reliability, reducing downtime, and managing operational risk in next-generation cloud-native ecosystems [dang2019aiops].
Existing research on automated IT operations generally falls into three major paradigms: rule-based expert systems, machine learning-based anomaly detection frameworks, and multi-agent automation architectures. Early expert systems [xu2017belief, muslewski2022expert, sarazin2021expert] offered interpretable and deterministic decision rules but required extensive manual knowledge engineering and were often brittle in the face of evolving infrastructure topologies. Machine learning–based approaches have achieved considerable success in anomaly detection and log pattern analysis; however, they typically lack contextual reasoning capabilities needed for accurate root cause localization and autonomous remediation [du2017deeplog, abid2021review, sangaiah2022automatic, he2021survey]. More recently, multi-agent systems (MAS) have been explored as a promising direction to enable distributed coordination, local reasoning, and decentralized decision-making in large-scale IT environments [wooldridge2009introduction, seitz2021automation, yin2025floorplan, wang2025enhancing, zhang2024adagent, zhang2025ccma, xu2024implementing, d2021designing]. Nevertheless, current MAS-based designs often suffer from several limitations, including inefficient information sharing, limited long-term memory, and poor adaptability to dynamic runtime conditions, particularly in highly elastic and heterogeneous cloud operations [d2021designing, pulikottil2021multi, li2024survey].
These limitations highlight the need for a new generation of intelligent, adaptive, and context-aware autonomous operation systems capable of perceiving complex system states, reasoning about causal structures, and coordinating actions across distributed environments[kephart2003vision, lake2017build, yuan2019causal, yao2021survey].
@@
Despite notable progress, current solutions face several critical limitations: (1) Information Overload: The explosive growth of operational data overwhelms existing analysis pipelines, degrading decision accuracy and response latency [he2021survey, ryciak2022anomaly]; (2) Poor Task Coordination: Lack of intelligent task decomposition and scheduling strategies leads to inefficient resource utilization [seitz2021automation, xu2024implementing]; (3) Context Loss: Traditional approaches fail to maintain crucial contextual information during long-running operations, impairing diagnosis continuity and repair consistency [liu2024lost, hu2025hiagent]; (4) Limited Adaptability: Most models fail to evolve with system drift or new failure modes, limiting their ability to generalize to unseen operational conditions [wellsandt2022hybrid, chen2022explainable].
@@
To address these challenges, this paper presents AOI (AI-Oriented Operations), a novel multi-agent collaborative framework designed for intelligent, self-adaptive IT operations management (Figure 2). AOI integrates specialized intelligent agents with large language model (LLM) reasoning and hierarchical memory to achieve scalable situational awareness and automated decision-making. Table I provides a comprehensive comparison of existing approaches with our proposed AOI framework, highlighting the unique combination of features that distinguish our work. The primary contributions of this work are summarized as follows:
@@
  • Multi-Agent Collaborative Architecture: We propose a specialized three-agent framework with clearly defined responsibilities-Observer for decision-making and coordination, Probe for safe information gathering, and Executor for system modifications, ensuring both safety and efficiency in operations.
@@
  • Intelligent Context Compression: We introduce an LLM-based Context Compressor with sliding-window mechanisms that selectively retains critical information while reducing context size by over 70%, enabling scalable operations in data-intensive environments.
@@
  • Dynamic Task Scheduling Strategy: We develop adaptive task scheduling algorithms that balances exploration (probing) and exploitation (execution) based on real-time system states and historical patterns, optimizing both MTTR and repair success rate.
@@
  • Three-Layer Memory Architecture: We design a hierarchical memory management system with raw context, task queues, and compressed context layers, enabling efficient information retrieval and cross-task reuse across multi-incident operations.
The remainder of this paper is organized as follows: Section II reviews related work in traditional automated IT operations and multi-agent systems for automated IT operations. Section III presents our AOI framework architecture and core algorithms. Section IV describes our experimental setup and evaluation methods. Section V presents comprehensive experimental results and analyses. Section VI discusses implications and limitations, and Section VII concludes with future research directions.
Figure 2: AOI Multi-Agent Collaborative Framework Architecture
II Related Work
Research related to automated IT operations spans traditional expert systems, machine learning–driven AIOps techniques, multi-agent architectures for distributed decision-making, and emerging advances in large language model (LLM) context management [joy2024aiops, openai2023gpt4]. This section reviews the evolution of these areas and highlights their limitations in addressing the challenges of modern cloud-native operational environments [soldani2018pains].
II-A Traditional Automated IT Operations
Early efforts toward automating IT operations primarily relied on rule-based expert systems and deterministic diagnostic engines. Classical expert systems such as belief rule-based reasoning frameworks demonstrated effectiveness in structured industrial environments [xu2017belief]. Similarly, knowledge-graph–driven expert systems have been applied to condition-based maintenance and operational optimization in industrial systems [sarazin2021expert, muslewski2022expert]. These systems offered transparent and interpretable decision-making but required extensive manual rule engineering and were limited in their ability to adapt to evolving infrastructures.
With the rise of DevOps practices, automation workflows increasingly incorporated Infrastructure as Code (IaC) and configuration management tools. Systematic analyses of IaC research have documented significant progress in automated provisioning, configuration reproducibility, and system deployment [rahman2019systematic]. Complementary surveys on DevOps practices highlight improvements in continuous delivery pipelines and operational efficiency through automated orchestration [leite2019survey]. However, despite these advances, IaC-based automation still lacks the intelligence required for autonomous fault diagnosis, contextual decision-making, and dynamic adaptation in large-scale distributed systems. Recent work on AI co-pilots for infrastructure management underscores this gap, advocating for more proactive and reasoning-capable assistants [nelavelli2025devops].
II-B Machine Learning in AIOps
Machine learning–based approaches, collectively known as AIOps, have emerged to overcome the limitations of static rule-based systems. A number of surveys provide comprehensive analyses of AIOps methods and real-world operational challenges [dang2019aiops, yeruva2023aiops, diaz2023joint, zhang2025survey]. Among these techniques, log-based anomaly detection has been a major focus. Deep learning models such as LogAnomaly [meng2019loganomaly] and earlier sequence-learning approaches such as DeepLog [du2017deeplog] have achieved impressive results in detecting operational anomalies from large-scale unstructured logs. These methods typically learn normal system behavior from historical data and flag statistically or semantically unusual patterns as potential incidents, offering a scalable alternative to manual log inspection.
Other branches of AIOps research focus on predictive analytics and resource management. Time-series forecasting models, such as recurrent neural networks applied to cloud workload prediction [zi2024time], have been used to support proactive scaling and performance optimization. Unsupervised and explainable fault diagnosis methods further enhance interpretability in dynamic systems [chen2022explainable]. Despite these advances, existing machine learning approaches typically operate in isolation—handling logs, metrics, or traces separately. They lack a holistic, context-aware understanding of system states, and most models fail to integrate reasoning, memory, and action planning capabilities necessary for truly autonomous IT operations [abid2021review, he2021survey]. As a result, they often generate high false-positive rates or require significant human intervention to validate and act on alerts, limiting their effectiveness in fast-paced, large-scale environments [hsieh2024comprehensivexai].
II-C Multi-Agent Systems for Automated IT Operations
Multi-agent systems (MAS) provide another promising paradigm for distributed automation. Foundational MAS literature highlights their advantages in decentralization, autonomy, and collaborative problem-solving [wooldridge2009introduction]. In operational environments, MAS architectures have been applied to distributed monitoring and anomaly detection, allowing agents to independently observe and report on system states across heterogeneous infrastructures [khosravifar2018anomaly]. Other works explore collaborative fault diagnosis using agent coordination strategies, enabling distributed decision-making and improved fault localization [shao2021novel, xu2023graph].
More recent research extends MAS to cloud-native and manufacturing environments. D’Aniello et al. demonstrate distributed operational management using cooperative agents in cloud-manufacturing scenarios [d2021designing], while adaptive resource management studies show how MAS can respond to dynamic workloads [seitz2021automation]. Xu et al. further illustrate MAS applications in autonomous supply chains, emphasizing task decomposition and coordination under uncertainty [xu2024implementing]. However, existing MAS designs often suffer from limited global context awareness, inefficient knowledge sharing, and difficulties maintaining long-term memory—constraints that hinder their ability to manage large-scale, highly dynamic IT infrastructures [pulikottil2021multi, li2024survey].
II-D Context Management and Compression in Large Language Models
Recent advances in large language models have introduced new opportunities for intelligent operations automation, but also new challenges related to long-context management [li2024surveyingmllm]. Long-document Transformer architectures such as Longformer [beltagy2020longformer], Reformer [kitaev2020reformer], and BigBird [zaheer2020big] provide scalable attention mechanisms to extend context windows [song2025transformersurvey]. However, they still face limitations when applied to continuously evolving, high-volume operational data streams.
To address context overload, several context compression and sparsification techniques have been proposed. LLMLingua [jiang2023llmlingua] compresses prompts to accelerate inference while preserving essential semantics, whereas contextual sparsity approaches such as DejaVu prune redundant information during inference [liu2023deja]. Complementary memory-augmented models, such as the Recurrent Memory Transformer (RMT) [bulatov2022recurrent], introduce persistent external memory to support long-range reasoning. Recent studies further reveal that LLMs exhibit position-dependent performance degradation in long contexts ("lost in the middle") [liu2024lost], and struggle with in-context exploration without explicit guidance [krishnamurthy2024can, peng2024securingllm].
While these methods improve efficiency and scalability, they are not designed specifically for IT operations environments, where context importance varies significantly across logs, metrics, events, and incidents [beyer2016sre]. Existing LLM context mechanisms lack adaptive prioritization and domain-specific reasoning capabilities needed for complex operational decision-making [lewis2020retrieval, borji2023categorical, liu2025comprehensive]. Emerging hierarchical memory architectures like HiAgent [hu2025hiagent] and retrieval-augmented in-context learning [chen2024retrieval] offer promising directions, yet remain unexplored in AIOps settings [syed2024ai, diaz2023joint].
TABLE I: Comparison of Related Work in Automated IT Operations
@@ II-D Context Management and Compression in Large Language Models
Approach Multi-Agent Context Mgmt Dynamic Scheduling LLM Integration Safety Guarantees
Rule-based Systems
Traditional AIOps Limited Limited
Agent-based Automation Limited Limited Limited
LLM-based Operations
Our AOI Framework
III Methodology
III-A Problem statement and framework overview

### v2

# AOI: Context-Aware Multi-Agent Operations via Dynamic Scheduling and Hierarchical Memory Compression

@@ Abstract
The proliferation of cloud-native architectures, characterized by microservices and dynamic orchestration, has rendered modern IT infrastructures exceedingly complex and volatile. This complexity generates overwhelming volumes of operational data, leading to critical bottlenecks in conventional systems: inefficient information processing, poor task coordination, and loss of contextual continuity during fault diagnosis and remediation. To address these challenges, we propose AOI (AI-Oriented Operations), a novel multi-agent collaborative framework that integrates three specialized agents with an LLM-based Context Compressor. Its core innovations include: (1) a dynamic task scheduling strategy that adaptively prioritizes operations based on real-time system states, (2) a three-layer memory architecture comprising Working, Episodic, and Semantic layers that optimizes context retention and retrieval. Extensive experiments on synthetic and real-world benchmarks show that AOI achieves 72.4% context compression while preserving 92.8% critical information, improves task success to 94.2%, and reduces MTTR by 34.4% over the best baseline. This work presents a paradigm shift towards scalable, adaptive, and context-aware autonomous operations, enabling robust management of next-generation IT infrastructures with minimal human intervention.
AOI: Context-Aware Multi-Agent Operations via Dynamic Scheduling and Hierarchical Memory Compression
Zishan Bai1, Jing Luo2, Ziyi Ni3, Enze Ge4, Jiacheng Shi5, Yichao Zhang6, Jiayi Gu7, Zhimo Han8, Riyang Bao9, Junfeng Hao10, 1Columbia University  2Chongqing University 3Institute of Automation, Chinese Academy of Sciences 4AI Agent Lab, Vokram Group  5College of William and Mary 6University of Texas  7Central University of Finance and Economics 8Zhengzhou University of Light Industry  9Emory University 10Department of Nephrology, Affiliated Hospital of Guangdong Medical University Correspondence: ygzhjf85@gmail.com
1 Introduction
@@
Modern IT infrastructures have evolved into highly distributed, cloud-native systems composed of microservices, containerized workloads, and dynamically orchestrated computing resources Dragoni et al. (2017); Yu (2025); Bao et al. (2026). While this architectural paradigm significantly enhances scalability and deployment agility, it also introduces unprecedented complexity and volatility into system behaviors Gannon et al. (2017). This shift not only surpasses the cognitive limits of human operators but also fundamentally challenges the design principles of automated operation systems, which must now reason about highly dynamic, interdependent, and large-scale system states Joy et al. (2024). As organizations scale their digital services, the volume of logs, monitoring metrics, and alerts grows exponentially, quickly surpassing the cognitive and operational limits of human operators Beyer et al. (2016). This challenge is well illustrated by large-scale production environments, where even a single misconfiguration or latent performance anomaly can propagate unpredictably across microservices, complicating diagnosis and dramatically increasing mean time to recovery (MTTR) Taibi et al. (2019). Empirical studies show that log data in modern distributed systems can reach terabyte-level daily volumes, making timely incident detection and remediation infeasible without automation Xu et al. (2009). Consequently, the demand for intelligent, autonomous IT operations systems has become critical for sustaining reliability, reducing downtime, and managing operational risk in next-generation cloud-native ecosystems Dang et al. (2019).
@@
Existing research on automated IT operations generally falls into three major paradigms: rule-based expert systems, machine learning-based anomaly detection frameworks, and multi-agent automation architectures. Early expert systems Xu et al. (2020); Muślewski et al. (2022); Sarazin et al. (2021); Yu et al. (2025b, c) offered interpretable and deterministic decision rules but required extensive manual knowledge engineering and were often brittle in the face of evolving infrastructure topologies. Machine learning–based approaches have achieved considerable success in anomaly detection and log pattern analysis; however, they typically lack contextual reasoning capabilities needed for accurate root cause localization and autonomous remediation Du et al. (2017); Abid et al. (2021); Sangaiah et al. (2022); He et al. (2021). More recently, multi-agent systems (MAS) have been explored as a promising direction to enable distributed coordination, local reasoning, and decentralized decision-making in large-scale IT environments Wooldridge (2009); Seitz et al. (2021); Yin et al. (2025); Wang et al. (2025); Zhang et al. (2024, 2025b); Xu et al. (2024); D’Aniello et al. (2021). Nevertheless, current MAS-based designs often suffer from several limitations, including inefficient information sharing, limited long-term memory, and poor adaptability to dynamic runtime conditions, particularly in highly elastic and heterogeneous cloud operations D’Aniello et al. (2021); Pulikottil et al. (2021); Li et al. (2024b). These limitations highlight the need for a new generation of intelligent, adaptive, and context-aware autonomous operation systems capable of perceiving complex system states, reasoning about causal structures, and coordinating actions across distributed environments Kephart and Chess (2003); Lake et al. (2017); Pearl et al. (2016); Yao et al. (2021); Ni et al. (2025a).
@@
Despite progress, existing solutions face key limitations: (1) Information Overload: Rapid growth of operational data overwhelms analysis pipelines, degrading decision accuracy and response latency He et al. (2021); Ryciak et al. (2022); (2) Poor Task Coordination: Lack of intelligent task decomposition and scheduling leads to inefficient resource utilization Seitz et al. (2021); Xu et al. (2024); (3) Context Loss: Traditional approaches fail to preserve crucial context in long-running operations, impairing diagnosis continuity and repair consistency Liu et al. (2024); Hu et al. (2025); (4) Limited Adaptability: Most models cannot evolve with system drift or new failure modes, limiting generalization to unseen conditions Wellsandt et al. (2022); Chen et al. (2022); Ni et al. (2025b).
@@
To address these challenges, we present AOI (AI-Oriented Operations), a novel multi-agent framework for intelligent self-adaptive IT operations management (Figure 1). AOI integrates specialized agents with LLM reasoning and hierarchical memory to enable scalable situational awareness and automated decision-making. Table 1 compares AOI with existing approaches and highlights its distinctive features. Our main contributions are:
Figure 2: AOI Multi-Agent Collaborative Framework Architecture. The framework consists of three specialized agents (Observer, Probe, Executor) coordinated through a central Context Compressor, with a three-layer memory architecture for efficient information management.
@@
  • Multi-Agent Collaborative Architecture: We propose a three-agent framework—Observer (coordination), Probe (safe information gathering), and Executor (system modification)—to ensure safe and efficient operations.
@@
  • Intelligent Context Compression: We introduce an LLM-based context compressor with sliding windows that preserves critical information while reducing context size by over 70
@@
  • Dynamic Task Scheduling: We develop adaptive scheduling that balances probing and execution based on real-time states and historical patterns to optimize MTTR and repair success.
@@
  • Three-Layer Memory Architecture: We design a hierarchical memory with raw context, task queues, and compressed context layers for efficient retrieval and cross-task reuse.
Table 1: Comprehensive Comparison of Related Work in Automated IT Operations
@@
Approach Multi-Agent Context Dynamic LLM Safety Memory Theoretical
Coordination Compression Scheduling Integration Guarantees Hierarchy Foundation
Expert Systems Xu et al. (2020); Sarazin et al. (2021)
DeepLog Du et al. (2017) Limited
LogAnomaly Meng et al. (2019) Limited
Traditional AIOps Dang et al. (2019) Limited Limited
MAS Automation Seitz et al. (2021) Limited Limited
Cloud-MAS D’Aniello et al. (2021) Limited Limited Limited
LLMLingua Jiang et al. (2023) Limited
HiAgent Hu et al. (2025) Limited
AOI (Ours)
2 Related Work
Research related to automated IT operations spans traditional expert systems, machine learning–driven AIOps techniques, multi-agent architectures for distributed decision-making, and emerging advances in large language model (LLM) context management Joy et al. (2024); OpenAI (2023). This section reviews the evolution of these areas and highlights their limitations in addressing the challenges of modern cloud-native operational environments Soldani et al. (2018).
2.1 Traditional Automated IT Operations
Early efforts toward automating IT operations primarily relied on rule-based expert systems and deterministic diagnostic engines. Classical expert systems such as belief rule-based reasoning frameworks demonstrated effectiveness in structured industrial environments Xu et al. (2020). Similarly, knowledge-graph–driven expert systems have been applied to condition-based maintenance and operational optimization in industrial systems Sarazin et al. (2021); Muślewski et al. (2022). These systems offered transparent and interpretable decision-making but required extensive manual rule engineering and were limited in their ability to adapt to evolving infrastructures.
Beyond rule-based expert systems, a line of research has investigated formal reliability and diagnosability properties of complex networked systems, providing theoretical foundations for fault detection and isolation. Prior studies have analyzed conditional diagnosability, connectivity, and fault tolerance of various interconnection networks under rigorous diagnostic models, demonstrating how structural properties directly impact system robustness and recovery guarantees. Such works offer valuable insights into how system topology and fault models influence safe operation and diagnosis, which motivates the safety-aware and reliability-driven design principles adopted in autonomous IT operations frameworks Wang et al. (2017b); Wang and Wang (2021); Wang et al. (2024, 2017a).
With the rise of DevOps practices, automation workflows increasingly incorporated Infrastructure as Code (IaC) and configuration management tools. Systematic analyses of IaC research have documented significant progress in automated provisioning, configuration reproducibility, and system deployment Rahman et al. (2019). Complementary surveys on DevOps practices highlight improvements in continuous delivery pipelines and operational efficiency through automated orchestration Leite et al. (2019). However, despite these advances, IaC-based automation still lacks the intelligence required for autonomous fault diagnosis, contextual decision-making, and dynamic adaptation in large-scale distributed systems. Recent work on AI co-pilots for infrastructure management underscores this gap, advocating for more proactive and reasoning-capable assistants Yang et al. (2025).
2.2 Machine Learning in AIOps
Machine learning–based approaches, collectively known as AIOps, have emerged to overcome the limitations of static rule-based systems. A number of surveys provide comprehensive analyses of AIOps methods and real-world operational challenges Dang et al. (2019); Yeruva and Ramu (2023); Diaz-De-Arcaya et al. (2023); Zhang et al. (2025a). Among these techniques, log-based anomaly detection has been a major focus. Deep learning models such as LogAnomaly Meng et al. (2019) and earlier sequence-learning approaches such as DeepLog Du et al. (2017) have achieved impressive results in detecting operational anomalies from large-scale unstructured logs. These methods typically learn normal system behavior from historical data and flag statistically or semantically unusual patterns as potential incidents, offering a scalable alternative to manual log inspection.
Other branches of AIOps research focus on predictive analytics and resource management. Time-series forecasting models, such as recurrent neural networks applied to cloud workload prediction Zi (2024), have been used to support proactive scaling and performance optimization. Unsupervised and explainable fault diagnosis methods further enhance interpretability in dynamic systems Chen et al. (2022). Despite these advances, existing machine learning approaches typically operate in isolation—handling logs, metrics, or traces separately. They lack a holistic, context-aware understanding of system states, and most models fail to integrate reasoning, memory, and action planning capabilities necessary for truly autonomous IT operations Abid et al. (2021); He et al. (2021). As a result, they often generate high false-positive rates or require significant human intervention to validate and act on alerts, limiting their effectiveness in fast-paced, large-scale environments Hsieh et al. (2024).
2.3 Multi-Agent Systems for Automated IT Operations
Multi-agent systems (MAS) provide another promising paradigm for distributed automation. Foundational MAS literature highlights their advantages in decentralization, autonomy, and collaborative problem-solving Wooldridge (2009). In operational environments, MAS architectures have been applied to distributed monitoring and anomaly detection, allowing agents to independently observe and report on system states across heterogeneous infrastructures Khosravifar (2018). Other works explore collaborative fault diagnosis using agent coordination strategies, enabling distributed decision-making and improved fault localization Shao et al. (2021); Xu et al. (2023).
More recent research extends MAS to cloud-native and manufacturing environments. D’Aniello et al. demonstrate distributed operational management using cooperative agents in cloud-manufacturing scenarios D’Aniello et al. (2021), while adaptive resource management studies show how MAS can respond to dynamic workloads Seitz et al. (2021). Xu et al. further illustrate MAS applications in autonomous supply chains, emphasizing task decomposition and coordination under uncertainty Xu et al. (2024). However, existing MAS designs often suffer from limited global context awareness, inefficient knowledge sharing, and difficulties maintaining long-term memory—constraints that hinder their ability to manage large-scale, highly dynamic IT infrastructures Pulikottil et al. (2021); Li et al. (2024b).
More broadly, complex real-world systems across different domains are often modeled as large-scale graphs, where robustness, efficiency, and structural properties play a central role in system behavior. Recent studies on multilayer transportation networks, geophysical systems, and graph-based learning further highlight the importance of structure-aware modeling and robustness analysis in complex environments, reinforcing the relevance of graph-structured abstractions in intelligent operational decision-making Li et al. (2022); Jiang et al. (2024); Pan et al. (2024).
2.4 Context Management and Compression in Large Language Models
Recent advances in large language models have introduced new opportunities for intelligent operations automation, but also new challenges related to long-context management Li et al. (2024a). Long-document Transformer architectures such as Longformer Beltagy et al. (2020), Reformer Kitaev et al. (2020), and BigBird Zaheer et al. (2020) provide scalable attention mechanisms to extend context windows Song et al. (2025). However, they still face limitations when applied to continuously evolving, high-volume operational data streams.
To address context overload, several context compression and sparsification techniques have been proposed. LLMLingua Jiang et al. (2023) compresses prompts to accelerate inference while preserving essential semantics, whereas contextual sparsity approaches such as DejaVu prune redundant information during inference Liu et al. (2023). Complementary memory-augmented models, such as the Recurrent Memory Transformer (RMT) Bulatov et al. (2022), introduce persistent external memory to support long-range reasoning. Recent studies further reveal that LLMs exhibit position-dependent performance degradation in long contexts ("lost in the middle") Liu et al. (2024); Ni et al. (2024), and struggle with in-context exploration without explicit guidance Krishnamurthy et al. (2024); Peng et al. (2024); Ni et al. (2025c).
While these methods improve efficiency and scalability, they are not designed specifically for IT operations environments, where context importance varies significantly across logs, metrics, events, and incidents Beyer et al. (2016). Existing LLM context mechanisms lack adaptive prioritization and domain-specific reasoning capabilities needed for complex operational decision-making Lewis et al. (2020); Borji (2023); Liu et al. (2025). Emerging hierarchical memory architectures like HiAgent Hu et al. (2025) and retrieval-augmented in-context learning Chen et al. (2024); Yu et al. (2025a) offer promising directions, yet remain unexplored in AIOps settings Syed and Anazagasty (2024); Diaz-De-Arcaya et al. (2023).
3 Theoretical Foundations
We establish the theoretical foundations of the AOI framework through formal analysis of multi-agent coordination, dynamic scheduling convergence, and context compression stability.
3.1 Markov Decision Process Formalization
We model the AOI multi-agent coordination as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP), defined by the tuple $\mathcal{M}=\langle\mathcal{N},\mathcal{S},\{\mathcal{A}_{i}\},\mathcal{T},\mathcal{R},\{\Omega_{i}\},\mathcal{O},\gamma\rangle$, where:
  • $\mathcal{N}=\{Observer,Probe,Executor\}$ is the set of agents
  • $\mathcal{S}$ is the joint state space of the IT infrastructure
  • $\mathcal{A}_{i}$ is the action space for agent $i$
  • $\mathcal{T}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is the state transition function
  • $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the shared reward function
  • $\Omega_{i}$ is the observation space for agent $i$
  • $\mathcal{O}:\mathcal{S}\times\mathcal{A}\times\Omega\rightarrow[0,1]$ is the observation function
  • $\gamma\in[0,1)$ is the discount factor
The optimal joint policy $\pi^{*}=\{\pi_{i}^{*}\}_{i\in\mathcal{N}}$ maximizes the expected cumulative discounted reward:
$\pi^{*}=\arg\max_{\pi}\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(s_{t},a_{t})\mid\pi\right]$ (1)
3.2 Dynamic Scheduling Convergence Analysis
We prove the convergence of our dynamic scheduling strategy using potential function analysis. Define the scheduling state at time $t$ as $\sigma_{t}=(q_{t},c_{t},h_{t})$, where $q_{t}$ is the task queue state, $c_{t}$ is the compressed context, and $h_{t}$ is the historical pattern embedding.
Theorem 1 (Scheduling Convergence).
Under the AOI dynamic scheduling strategy with exploration-exploitation parameter $\lambda\in(0,1)$, the expected task completion time converges to within a factor of $(1+\epsilon)$ of the optimal static scheduling policy for any $\epsilon>0$, with probability at least $1-\delta$, provided the number of scheduling iterations $T\geq\frac{|\mathcal{A}|\log(|\mathcal{A}|/\delta)}{\epsilon^{2}\lambda^{2}}$.
Proof.
Define the potential function $\Phi(\sigma_{t})=\sum_{\tau\in q_{t}}w(\tau)\cdot d(\tau,t)$, where $w(\tau)$ is the priority weight and $d(\tau,t)$ is the delay cost. Under our scheduling policy:
$\mathbb{E}[\Phi(\sigma_{t+1})-\Phi(\sigma_{t})\mid\sigma_{t}]\leq-\eta\cdot\|\nabla_{\sigma}\Phi(\sigma_{t})\|^{2}+\frac{\lambda^{2}L^{2}}{2}$ (2)
where $\eta>0$ is the learning rate and $L$ is the Lipschitz constant of the reward function. By applying the Robbins-Monro stochastic approximation theorem, the scheduling policy converges to a stationary point. Combined with the UCB-style exploration bonus, regret analysis yields the stated bound. ∎
3.3 Context Compression Stability
We analyze the stability of the LLM-based context compression through Lyapunov analysis. Let $\mathcal{C}(t)$ denote the context state at time $t$, with compression operator $\mathcal{F}:\mathcal{C}\rightarrow\mathcal{C}_{compressed}$.
Theorem 2 (Compression Stability).
The context compression mechanism with sliding window size $w$ and overlap ratio $\rho\in(0.5,1)$ satisfies Lyapunov stability, i.e., there exists a Lyapunov function $V:\mathcal{C}\rightarrow\mathbb{R}_{\geq 0}$ such that:
$V(\mathcal{F}(\mathcal{C}))-V(\mathcal{C})\leq-\kappa\cdot\|\mathcal{C}-\mathcal{C}^{*}\|^{2}$ (3)
for some $\kappa>0$ and equilibrium state $\mathcal{C}^{*}$, ensuring bounded information loss.
Proof.
Define $V(\mathcal{C})=\frac{1}{2}\|\mathcal{C}-\mathcal{C}^{*}\|_{\mathcal{H}}^{2}$, where $\|\cdot\|_{\mathcal{H}}$ is the norm in the reproducing kernel Hilbert space induced by the LLM embedding. The compression operator satisfies:
$\|\mathcal{F}(\mathcal{C})-\mathcal{C}^{*}\|_{\mathcal{H}}\leq(1-\rho)\|\mathcal{C}-\mathcal{C}^{*}\|_{\mathcal{H}}+\epsilon_{LLM}$ (4)
where $\epsilon_{LLM}$ is the LLM approximation error bounded by model capacity. For $\rho>0.5$, contraction holds and stability follows. ∎
3.4 Information Preservation Bound
We establish theoretical guarantees on information preservation during compression.
Theorem 3 (Information Preservation).
Let $I(\mathcal{C};\mathcal{Y})$ denote the mutual information between context $\mathcal{C}$ and operational outcome $\mathcal{Y}$. The compressed context $\mathcal{C}_{comp}=\mathcal{F}(\mathcal{C})$ satisfies:
$I(\mathcal{C}_{comp};\mathcal{Y})\geq(1-\epsilon_{info})\cdot I(\mathcal{C};\mathcal{Y})$ (5)
where $\epsilon_{info}=O\left(\frac{1}{w\cdot\rho}\right)$ depends on window size $w$ and overlap ratio $\rho$.
This bound theoretically justifies our empirical observation that larger window sizes improve information preservation (Table 4).
4 Methodology
4.1 Problem statement and framework overview